OpenStudy (nuccioreggie):

A kilobyte is 210 bytes and a megabyte is 220 bytes. How many kilobytes are in a megabyte? 2200 kilobytes 230 kilobytes 210 kilobytes 2 kilobytes

OpenStudy (nuccioreggie):

@AnswerMyQuestions @jhonyy9

OpenStudy (anonymous):

\[\frac{1\text{ kb}}{210\text{ b}}\times\frac{220\text{ b}}{1\text{ mb}}\approx\frac{1.04\text{ kb}}{1\text{ mb}}\] Are you sure you have those conversion rates right?

OpenStudy (nuccioreggie):

yea whats the answer